# Electronic reconstruction and interface engineering of emergent spin fluctuations in compressively strained La3Ni2O7 on SrLaAlO4(001)

### v1

@@ Abstract
Motivated by the recent observation of ambient-pressure superconductivity with K in La3Ni2O7 on SrLaAlO4(001), we explore the structural and electronic properties as well as the spin-spin correlation function of this bilayer nickelate system by using density functional theory including a Coulomb repulsion term. We find that the compressive strain exerted by this substrate leads to an unconventional occupation of the antibonding Ni states, distinct from the superconducting bulk compound under pressure. While pure strain effects rather modestly enhance the dynamical spin susceptibility via Fermi surface nesting, investigation of a reconstructed interface composition as observed in transmission electron microscopy uncovers a strong amplification of the spin fluctuations. These results provide insights into the emergence of superconductivity in strained La3Ni2O7, suggest a possible key role of the interface, and highlight fundamental differences from the hydrostatic pressure scenario.
@@ I Introduction
Here we explore the structural and electronic properties of La3Ni2O7/SrLaAlO4(001) by using density functional theory including a Coulomb repulsion term, with an explicit treatment of the interface. In addition to the ideal interface stacking, we consider a reconstructed interface composition that has recently been identified in transmission electron microscopy (TEM) [53]. Structural analysis reveals close agreement with recent x-ray diffraction (XRD) and TEM data for La3Ni2O7 on different prominent substrates [57] and emphasizes the importance of the interfacial reconstruction. We observe finite octahedral tilts throughout the bilayer nickelate, which surprisingly are quenched at the reconstructed interface due to geometric constraints. We find that compressive strain in the presence of a substrate induces an electronic reconstruction in La3Ni2O7, characterized by an unconventional partial occupation of the antibonding Ni states. Interestingly, while strain alone leads to a rather modest enhancement of the dynamical spin susceptibility as a result of Fermi surface nesting, the reconstructed interface composition substantially amplifies the spin fluctuations. A detailed disentanglement of the scattering channels reveals that the amplified fluctuations are driven by the emergent interfacial Ni states.
@@ II.1 Interface geometry and electronic reconstruction in La3Ni2O7/SrLaAlO4(001)
Figure 2 shows the projected momentum-resolved density of states, , where denotes Ni manifolds at different sites . In general, compressive strain drives a significant charge transfer from the Ni reservoir to the Ni states, since the concomitant vertical expansion of the nickelate region tends to lower the energy of the Ni orbital relative to Ni . Intriguingly, this allows to control the Ni orbital polarization over a considerably larger interval than hydrostatic pressure [50]. Here, we find that this mechanism leads to an unconventional occupation of the antibonding Ni states around the point [Fig 2(b)], which are empty in the bulk [Fig 2(a)], even at finite pressure. Their occupation is fairly uniform among the different layers, signaling the absence of electrostatic doping. Simultaneously, the bonding Ni states are entirely filled and their flat-band-like maximum is lowered to eV in all layers.
@@
Importantly, this picture applies to the system with reconstructed interface as well [Fig 2(c)]. Due to the Al substitution in layer L1, the adjacent layer L2 presents a single Ni band instead of a bonding-antibonding pair [Fig 2(c)]. It is more strongly filled than the other Ni bands in either the ideal or the reconstructed system and provides a substantial contribution to the total density of states at the Fermi energy [Fig. 1(b)], which plays a central role in the spin susceptibility enhancement (see below). Analysis of the layer polarities leaves the NiO2 layer L2 in a formal state, leading to an altered local Ni3+ valence similar to bulk LaNiO3. Interestingly, this interfacial NiO2 layer is embedded in a highly anisotropic chemical environment, which induces a considerable Ni orbital polarization. This effect is even more pronounced than in LaNiO3/LaAlO3(001) superlattices, which have been widely explored in the context of Fermi surface engineering and nickelate superconductivity [58, 59].
@@
The Fermi surface of La3Ni2O7 is known to consist of two characteristic sheets: an sheet of predominantly Ni character and a sheet with some admixture of Ni [4, 5, 6, 27]. Under hydrostatic pressure [1, 4, 5, 6, 7, 21, 16] or tensile strain as imposed by SrTiO3(001) [50], the system undergoes a topological transition, leading to the emergence of a hole pocket due to a metallization of the bonding Ni states. In stark contrast, our simulations for compressively strained La3Ni2O7/SLAO(001) reveal a distinct Fermi surface composition, reflecting the differences in electronic reconstruction [Fig. 2(d)]. The sheet exhibits an enhanced Ni contribution compared to the bulk, while the sheet appears more rectangular and features a significantly promoted Ni character. Additional electron pockets emerge near the point, originating from the partial occupation of the antibonding Ni states (). For the reconstructed interface composition, we find a notable splitting of the sheet. Moreover, a larger electron pocket around the point can be observed, predominantly associated with Ni states localized in layer L2, and a smaller pocket that stems from layer L3 [Fig. 2(c,d)]. These results highlight the effect of interface reconstructions on the electronic structure and suggest their critical role in the modulation of the superconducting properties. We will corroborate this aspect by calculating the dynamical spin susceptibility below.
@@ II.3 Spin susceptibility in La3Ni2O7/SrLaAlO4(001)
The superconducting pairing is determined by the dynamical spin susceptibility , which we compare in Fig. 4 considering a Kanamori-type interaction vertex with and eV in the random phase approximation (RPA). For completeness, we also show the bare Lindhard susceptibility .
The spin susceptibility of bulk La3Ni2O7 at ambient pressure is dominated by intraorbital scattering [50]. Direct comparison reveals that La3Ni2O7/SLAO(001) exhibits a moderately enhanced magnitude and sharpness of the characteristic susceptibility peaks along the direction. The difference can be traced back to the intraorbital and interorbital - channels (not shown).
Intriguingly, we find a very strong enhancement of the susceptibility for the La3Ni2O7/SLAO(001) system if we consider the reconstructed interface geometry observed in TEM. These changes are not only quantitative, but also qualitative, as they are driven by considerably increased intraorbital scattering and further promoted by contributions from the interorbital - channel (Fig. 4, bottom row). Both clearly surpass the channel.
We identify that this phenomenon originates predominantly from the emergent Ni states localized near the interface, which contribute strongly to the density of states at the Fermi level and exhibit favorable Fermi surface nesting properties (Figs. 1 and 2). Together, these effects result in the strong amplification of the spin fluctuations. Considering that recent experiments reporting ambient-pressure superconductivity in strained systems are performed on capped bilayer nickelate films of 4 to 6 bilayers thickness [53, 54, 55], these findings suggest that interfacial effects may play a key role in the emergence of superconductivity in La3Ni2O7 on SLAO(001).
@@ III Discussion
We investigated the effect of compressive strain and interface reconstructions in La3Ni2O7 on SrLaAlO4(001) by performing first-principles simulations including a Coulomb repulsion term, treating the interface explicitly.
Comparison of the structural optimization results, specifically the evolution of the vertical lattice parameter of the bilayer nickelate on different technologically relevant substrates, with recent x-ray diffraction and transmission electron microscopy measurements on strained epitaxial thin films revealed close agreement and highlighted the key role of the reconstructed interface composition. We observed finite octahedral tilts throughout the bilayer nickelate. Intriguingly, these tilts are suppressed at the reconstructed interface, accompanied by significantly enhanced spin fluctuations. This finding provides an important contribution to the ongoing debate on whether superconductivity in bilayer nickelates requires a structural transition to the high-symmetry phase or can arise from purely electronic effects [24, 50].
We found that compressive strain drives an electronic reconstruction in the bilayer nickelate, characterized by an unconventional occupation of the antibonding Ni states with simultaneous absence of interfacial charge transfer. This scenario contrasts sharply with the notion of metallized bonding Ni states, considered to be critical for the superconducting state in pressurized bulk La3Ni2O7. The concomitant substantial modifications of the Fermi surface and amplified spin fluctuations, especially for the reconstructed interface due to the emergent Ni states, point to an electron pairing that deviates fundamentally from the flat-band physics dominant under hydrostatic pressure.
@@
Acknowledgements.
This work was supported by the National Science Foundation, Grant No. NSF-DMR-2118718.IV Methods
@@ IV.1 Density functional theory calculations
We performed first-principles calculations in the framework of density functional theory [62] (DFT) as implemented in the Quantum Espresso code [63]. The generalized gradient approximation was used for the exchange and correlation functional as parametrized by Perdew, Burke, and Ernzerhof [64] in conjunction with ultrasoft pseudopotentials [65], as successfully employed in previous work [66, 67, 61]. Static correlation effects were considered within the DFT formalism [68], employing eV on the Ni and Ti states [69, 70, 71, 1, 16, 50]. Octahedral rotations are fully accounted for by using supercells with two transition metal sites per layer, setting to the substrate lattice parameter of [53, 57] and accurately optimizing as well as the ionic positions. The symmetric supercells contain two equivalent interfaces, 3 bilayers of La3Ni2O7, and 5 layers of substrate, corresponding to 142 atoms (Fig. 1). Wave functions and density were expanded into plane waves up to cutoff energies of and Ry, respectively. The Brillouin zone was sampled by using a Monkhorst-Pack -point grid [72] and mRy Methfessel-Paxton smearing [73]. The ionic positions were accurately optimized, reducing ionic forces below mRya.u. Subsequently, Fermi surfaces and densities of states were obtained by using -point grids. Reference calculations for ambient-pressure bulk La3Ni2O7 () were performed by using 48-atom supercells [24, 34] with and -point grids.

### v2

@@ Abstract
Motivated by the recent observation of ambient-pressure superconductivity with K in La3Ni2O7 on SrLaAlO4(001) [53, 55], we explore the structural and electronic properties as well as the spin-spin correlation function of this bilayer nickelate system by using density functional theory including a Coulomb repulsion term. We find that the compressive strain exerted by this substrate leads to an unconventional occupation of the antibonding Ni states around the point, distinct from the superconducting bulk compound under pressure. While pure strain effects rather modestly enhance the dynamical spin susceptibility, investigation of a reconstructed interface composition as observed in transmission electron microscopy uncovers a strong amplification of the spin fluctuations due to Fermi surface nesting of the antibonding Ni states near the interface. These results provide insights into the emergence of superconductivity in strained La3Ni2O7, suggest a possible key role of the interface, and highlight fundamental differences from the hydrostatic pressure scenario.
@@ I Introduction
Here we explore the structural and electronic properties of La3Ni2O7/SrLaAlO4(001) by using density functional theory including a Coulomb repulsion term, with an explicit treatment of the interface. In addition to the ideal interface stacking, we consider a reconstructed interface composition that has been identified in transmission electron microscopy (TEM) [53]. Structural analysis reveals close agreement with recent x-ray diffraction (XRD) and TEM data for La3Ni2O7 on different prominent substrates [57] and emphasizes the importance of the interfacial reconstruction. We observe finite octahedral tilts throughout the bilayer nickelate, which surprisingly are quenched at the reconstructed interface due to geometric constraints. We find that compressive strain in the presence of a substrate induces an electronic reconstruction in La3Ni2O7, characterized by an unconventional partial occupation of the antibonding Ni states. Interestingly, while compressive strain alone leads to a rather modest enhancement of the dynamical spin susceptibility, the reconstructed interface composition substantially amplifies the spin fluctuations. A detailed disentanglement of the scattering channels reveals that this phenomenon is driven by strong Fermi surface nesting involving the emergent interfacial Ni states.
@@ II.1 Interface geometry and electronic reconstruction in La3Ni2O7/SrLaAlO4(001)
Figure 2 shows the projected momentum-resolved density of states, , where denotes Ni manifolds in different layers . In general, compressive strain drives a significant charge transfer from the Ni reservoir to the Ni states, since the concomitant vertical expansion of the nickelate region tends to lower the energy of the Ni orbital relative to Ni . Intriguingly, this allows to control the Ni orbital polarization over a considerably larger interval than hydrostatic pressure [50]. Here, we find that this mechanism leads to an unconventional occupation of the antibonding Ni states around the point [Fig 2(b)], which are empty in the bulk [Fig 2(a)], even at finite pressure. Their occupation is fairly uniform throughout the different layers, signaling the absence of electrostatic doping. Simultaneously, the bonding Ni states are entirely filled and their flat-band-like maximum is universally lowered to eV.
@@
Importantly, this picture applies to the system with reconstructed interface as well [Fig 2(c)]. We observe the flat band now at eV in layer L3. Due to the Al substitution in layer L1, the adjacent layer L2 presents a single Ni band () instead of a bonding-antibonding pair. It is more strongly filled than the other Ni bands in either the ideal or the reconstructed system and provides a substantial contribution to the total density of states at the Fermi energy [Fig. 1(b)], which plays a central role in the enhancement of the spin susceptibility (see below). Analysis of the layer polarities leaves the NiO2 layer L2 in a formal state, leading to an altered local Ni3+ valence similar to bulk LaNiO3. Interestingly, this interfacial NiO2 layer is embedded in a highly anisotropic chemical environment, which induces a considerable Ni orbital polarization. This effect is even more pronounced than in LaNiO3/LaAlO3(001) superlattices, which have been widely explored in the context of Fermi surface engineering and nickelate superconductivity [58, 59].
@@
The Fermi surface of La3Ni2O7 is known to consist of two characteristic sheets: an sheet of predominantly Ni character and a sheet with some admixture of Ni [4, 5, 6, 27]. Under hydrostatic pressure [1, 4, 5, 6, 7, 21, 16] or tensile strain as imposed by SrTiO3(001) [50], the system undergoes a topological transition, i.e., the emergence of the hole pocket due to a metallization of the bonding Ni states. In stark contrast, our simulations for compressively strained La3Ni2O7/SLAO(001) reveal a distinct Fermi surface composition, reflecting the differences in electronic reconstruction [Fig. 2(d)]. The sheet exhibits an enhanced Ni contribution compared to the bulk, while the sheet appears more rectangular and features a significantly promoted Ni character. Most importantly, electron pockets emerge near the point, originating from the partial occupation of the antibonding Ni states. Specifically, for the reconstructed interface composition, we find a slight splitting of the sheet, a larger electron pocket around the point predominantly associated with Ni states localized in layer L2, and a smaller pocket that stems from layer L3 [Fig. 2(c,d)].
The pronounced effect of compressive strain and interface reconstructions on the electronic structure suggests their critical role in modulating the superconducting properties. We will corroborate this aspect by calculating the dynamical spin susceptibility below.
@@ II.3 Spin susceptibility in La3Ni2O7/SrLaAlO4(001)
The superconducting pairing is determined by the dynamical spin susceptibility , which we compare for different systems in Fig. 4(a) considering a Kanamori-type interaction vertex with and eV in the random phase approximation (RPA). For completeness, we also show the bare Lindhard susceptibility .
The spin susceptibility of bulk La3Ni2O7 at ambient pressure is dominated by intraorbital scattering [50]. Direct comparison reveals that La3Ni2O7/SLAO(001) exhibits a moderately enhanced magnitude and sharpness of the characteristic susceptibility peaks along the direction [Fig. 4(a)]. The difference can be traced back to the intraorbital and interorbital - channels (not shown).
Intriguingly, we find a very strong enhancement of the susceptibility for La3Ni2O7/SLAO(001) if we consider the reconstructed interface geometry observed in TEM [Fig. 4(a)]. These changes are not only quantitative, but also qualitative, as they are driven by considerably increased intraorbital scattering and further promoted by contributions from the interorbital - channel [Fig. 4(b)]. Both clearly surpass the channel.
This surprising enhancement of spin fluctuation weight originates predominantly from the emergent Ni states localized near the interface (i.e., the sheet), which contribute strongly to the density of states at the Fermi level (Figs. 1 and 2) and exhibit favorable nesting properties with the sheets, as indicated by the red vector in Fig. 4(c). Together, these effects result in the strong amplification of the spin fluctuations. Considering that recent experiments reporting ambient-pressure superconductivity in strained systems are performed on capped bilayer nickelate films of 4 to 6 bilayers thickness [53, 54, 55], these findings suggest that interfacial effects may play a key role in the emergence of superconductivity in La3Ni2O7 on SLAO(001).
@@ III Discussion
We investigated the effect of compressive strain and interface reconstructions in La3Ni2O7 on SrLaAlO4(001) by performing first-principles simulations including a Coulomb repulsion term and explicitly treating the interface. We found that compressive strain drives an electronic reconstruction in the bilayer nickelate, characterized by an unconventional occupation of the antibonding Ni states despite the absence of interfacial charge transfer. This scenario contrasts sharply with the notion of metallized bonding Ni states, considered to be critical for the superconducting state in pressurized bulk La3Ni2O7. The concomitant substantial modifications of the Fermi surface and amplified spin fluctuations, especially for the reconstructed interface due to strong nesting of the cuprate-shaped sheets with the emergent Ni states, point to an electron pairing that deviates fundamentally from the flat-band physics dominant under hydrostatic pressure, as illustrated in Fig. 4(c). We provide a tight-binding model (see Methods) that allows a detailed exploration of the gap structure and symmetry of the order parameter in future work.
Interestingly, recent angle-resolved photoemission spectroscopy (ARPES) experiments published after initial submission of this paper identified a suppressed contribution of the bonding Ni flat band (i.e., the hole pocket) to the Fermi surface [62, 63, 64], with the respective states shifted to at least meV below the Fermi level [63]. This behavior is in stark contrast to that of pressurized bulk samples and in qualitative agreement with our findings, as summarized in Fig. 4(c). Moreover, it has been suggested that conduction is localized at the interface [62], implying a layer dependence of the electronic structure throughout the sample, as also found by our simulations treating realistic interface compositions. Although the imaging resolution is limited so far, these findings raise a critical question: If the bonding Ni states are not active at the Fermi level, how does the electronic structure of compressively strained La3Ni2O7 differ from that at ambient pressure, and why does superconductivity emerge? Suggesting a tentative answer, we show that electron pockets formed by antibonding Ni states around the point may play a relevant role and that interface-related phenomena—such as Sr and Al interdiffusion into the bilayer stacking, as experimentally observed [53, 62]—possibly enhance superconductivity. Recent thin-film ARPES data provide indications that support this hypothesis [62, 63], as illustrated in Fig. 4(c), which calls for further investigation.
Comparison of the structural optimization results, specifically the evolution of the vertical lattice parameter of the bilayer nickelate on different technologically relevant substrates, with recent x-ray diffraction and transmission electron microscopy measurements on strained epitaxial thin films revealed close agreement and further corroborated the key role of the reconstructed interface composition. We observed finite octahedral tilts throughout the bilayer nickelate. Intriguingly, these tilts are suppressed at the reconstructed interface, accompanied by significantly enhanced spin fluctuations. This finding provides an important contribution to the ongoing debate on whether superconductivity in bilayer nickelates requires a structural transition to the high-symmetry phase or can arise from purely electronic effects [24, 50].
@@
IV Methods
@@ IV.1 Density functional theory calculations
We performed first-principles calculations in the framework of density functional theory [65] (DFT) as implemented in the Quantum Espresso code [66]. The generalized gradient approximation was used for the exchange and correlation functional as parametrized by Perdew, Burke, and Ernzerhof [67] in conjunction with ultrasoft pseudopotentials [68], as successfully employed in previous work [69, 70, 61]. Static correlation effects were considered within the DFT formalism [71], employing eV on the Ni and Ti states [72, 73, 74, 1, 16, 50]. Octahedral rotations are fully accounted for by using supercells with two transition metal sites per layer, setting to the substrate lattice parameter of [53, 57] and accurately optimizing as well as the ionic positions. The symmetric supercells contain two equivalent interfaces, 3 bilayers of La3Ni2O7, and 5 layers of substrate, corresponding to 142 atoms (Fig. 1). This setup ensures that key interfacial effects are fully captured. Additional inner layers would primarily reflect bulk-like behavior and would not alter the essential physics. Wave functions and density were expanded into plane waves up to cutoff energies of and Ry, respectively. The Brillouin zone was sampled by using a Monkhorst-Pack -point grid [75] and mRy Methfessel-Paxton smearing [76]. The ionic positions were accurately optimized, reducing ionic forces below mRya.u. Subsequently, Fermi surfaces and densities of states were obtained by using -point grids. Reference calculations for ambient-pressure bulk La3Ni2O7 () were performed by using 48-atom supercells [24, 34] with and -point grids.
@@ IV.3 Spin susceptibility calculations
Data availability
The data is available upon reasonable request to the authors.
Acknowledgements.
We thank Christine Ah-Yeung, Martin Bluschke, and Steef Smit for valuable discussions on the angle-resolved photoemission spectroscopy of bilayer nickelates. This work was supported by the National Science Foundation, Grant No. NSF-DMR-2118718.Author Contributions
BG and PJH conceived of the project. JJH, GRS, RGH, and PJH supervised the research. BG performed the theoretical simulations and corresponding analysis. BG and PJH wrote the paper. All authors discussed the results and revised the paper.
Competing Interests
The authors declare no competing interests.